OpenStudy (anonymous):

A population is normally distributed a. What % of the population lies to the left of the population z score -1.13 b.what % of the population lies to the left of the population z score 1.28?

OpenStudy (anonymous):

do you happen to know how to figure that out? I'm going to have questions similar on a test

OpenStudy (anonymous):

You'll probably be equipped with a Z table, right?

OpenStudy (anonymous):

no, its an online class. the exam wont have it on there. I'm going to see if I can shrink it enough to put onto my note card though

OpenStudy (anonymous):

Hold on, it's an online course? What's stopping you from using Wolfram to compute then? Anyway, I doubt you won't be given a table. The alternative would be to integrate a really hard function: \[\large P(Z<k)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^ke^{-x^2/2}~dx\] Now, there are two types of tables you could use, and you have to be careful about the type of table you're given. Left-tail tables (like this one http://dsearls.org/courses/M120Concepts/ClassNotes/Statistics/520A_LeftTailTable.htm) give you the proportion/probability of a population that resides to the left of the critical value \(k\), or \(P(Z<k)\):  Right-tail tables (like this one http://www.stat.cmu.edu/~rsteorts/btheory2/code/z-table.pdf) give the probability of \(P(Z>k)\):